1. If you invest $5,000 today at an annual interest rate of 5.35%, how much money will you have for your daughter's college education in 18 years?

$5,603

$10,000

$12,776

There is not enough information to answer this question.
Answer #1

Assuming annual compounding

We see that the amount acumulated is given as equal to=Present Value*(1+rate)^t=5000*(1+5.35%)^18=12775.906457528